# Momentor: Advancing Video Large Language Model with Fine-Grained Temporal Reasoning

### v1

@@ 1 Introduction
Inspired by the success of ChatGPT (OpenAI, 2022), numerous studies across various fields are attempting to integrate Large Language Models (LLMs) with their domain-specific tasks, seeking to bring innovation to these fields. For example, Video Large Language Models (Video-LLMs) such as VideoChat (Li et al., 2023b) and Video-ChatGPT (Maaz et al., 2023) adapt LLM to video modality, striving to merge the understanding, reasoning and interactive skills of LLM with video perception. They typically sample multiple frames from the video, use an image encoder to encode these frames separately, and employ a projection layer (e.g. a linear layer or Q-Former (Li et al., 2023a)) to adapt the visual features to the feature space of an open-source LLM ((Touvron et al., 2023), (Chiang et al., 2023)). By training on video-level captioning and QA tasks, they establish coarse-grained multimodal feature alignment and acquire the capability of instruction following.
@@
We conduct extensive experiments with our proposed Momentor. The results indicate that our Momentor outperforms previous Video-LLMs in multiple tasks involving precise temporal position, such as temporal grounding, dense captioning, action segmentation, and highlight detection. Momentor demonstrates advanced proficiency in temporal perception. It can provide appropriate responses based on user-indicated segments as well as quickly locate target segments that meet user requirements.
@@ 2 Related Work
2.1 Video Large Language Models
Many efforts have been made to transfer the task-handling capability of Large Language Models (LLMs) to the video modality, enabling them to complete various tasks based on video content in accordance with user instructions (Li et al., 2023b; Maaz et al., 2023; Zhang et al., 2023; Luo et al., 2023). Despite being effective in captioning or QA on short videos, the lack of fine-grained temporal modeling in these models prevents them from understanding or locating specific segments in long videos. In contrast, Momentor employs a Temporal Perception Module that integrates a continuous temporal token space for precise temporal positioning and modeling.
@@ 2.2 Temporally Grounded Video Understanding
Fine-grained video understanding tasks usually demand the model to view a video as a series of interconnected events and comprehend or locate them in a temporally grounded manner. For instance, action segmentation (Singh et al., 2016; Sarfraz et al., 2021; Du et al., 2022; Behrmann et al., 2022) requires the model to temporally split the video and output the action label for each segment; temporal grounding (Gao et al., 2017; Zhang et al., 2020b, a) demands the model to identify the start and end timestamps of the video segment corresponding to a given natural language query; highlight detection (Lei et al., 2021; Lin et al., 2023) requires the model to find out the central event in a video from a natural language description and pinpoint all related segments; dense video captioning (Alwassel et al., 2021; Yang et al., 2023) requires the model to list out all events contained in a video along with their start and end timestamps. Previous methods typically train a task-specific model for each task, whereas we aim to design a unified Video-LLM that can solve these tasks in a zero-shot manner.
@@ 3.2 Temporal Perception Module (TPM)
Since each sampled frame is encoded and projected separately, their features do not contain the corresponding temporal position information. After constructing a continuous temporal token space and applying the neighboring token propagation, now we can actually obtain temporal embeddings corresponding to any timestamp, which contain precise temporal position information and possess the valuable property of temporal continuity. Therefore, we obtain the temporal embeddings at the positions of the sampled frames and directly add them to the frame features, serving as a form of temporal position encoding to inject fine-grained temporal information.
@@ 3.3 Grounded Event-Sequence Modeling
After the stage of modality alignment, the model only learns the coarse-grained correspondence between visual and textual data. It still lacks fine-grained temporal awareness, so fine-tuning it directly on instruction data with precise timestamps can lead to slow convergence and ineffective event-sequence modeling. Therefore, we apply event-sequence decoding as an intermediary task that bridges the gap between low-level semantic alignment and high-level conceptual interaction. To be precise, given an untrimmed video as input, we require the model to output the event-sequence within it. We represent the $k$ th event as $E_{k}=[t_{start}^{k},t_{end}^{k},w_{1}^{k},...,w_{l_{k}}^{k}]$, where $t_{start}^{k}$, $t_{end}^{k}$ are the continuous temporal embeddings at the start and end of the $k$ th event, and $[w_{1}^{k},...,w_{l_{k}}^{k}]$ is a general caption composed of $l_{k}$ tokens for this event. The timestamps and general captions of each event in the event-sequence can be conveniently obtained during the construction of the instance-event matrix, without additional calculation (see Section 4.1). We concatenate all the events to formulate the event-sequence $T_{E}=\{E_{i}\}_{i=1}^{N_{E}}$, where $N_{E}$ is the number of events in the untrimmed video. We apply a language modeling loss for event-sequence decoding:
@@ 4.1 Structured Information Extraction
Model Temporal Grounding Highlight Detection
@@
Table 3: Comparison with existing Video-LLMs on temporal grounding and highlight detection
@@
Based on the result of instance tracking and event segmentation, we construct an Instance-Event Matrix, where each row represents an instance track (the video itself also counts as a track), and each column represents an event. Then we traverse the matrix and utilize several multimodal pre-trained models to extract visual clues such as scenes, instances, actions and attributes from each track. With the structured information organized in instance-event matrix, we can quickly generate instruction data that includes various spatio-temporal associations.
@@ 4.2 Instruction Generation
We feed the information in the instance-event matrix into Vicuna (Chiang et al., 2023), an open-source text-based LLM, to generate instruction data. We design various types of instruction-following tasks to comprehensively train and evaluate Video-LLMs. We incorporate 5 tasks focusing on single segment understanding as well as 3 tasks that involve reasoning across multiple segments, as shown in Figure 3. We utilize various prompts to guide Vicuna in generating instruction data for different tasks. Detailed task descriptions and prompts can be found in Appendix C and D. We select a substantial number of videos from YTTemporal-1B (Zellers et al., 2022) to build Moment-10M. Figure 4 shows the distribution of each type of instruction data in Moment-10M. As shown in Table 1, Moment-10M comprises 10 million instruction data over 1.5 million segments and 451.5 thousand instance tracks. On average, each video contains 22.7 segments, which reflects the complexity of the event-sequences in the videos. We fine-tune Momentor on Moment-10M, enabling it to perform segment-level reasoning and localization.
Figure 4: Distribution of different tasks in Moment-10M.
@@ 5.1 Experiment Setup
To comprehensively evaluate Momentor in fine-grained understanding and precise localization, we assess it in a zero-shot setting across four tasks, i.e., action segmentation, dense video captioning, temporal grounding, and highlight detection, using datasets such as Breakfast (Kuehne et al., 2014), 50 Salads (Stein & McKenna, 2013), ActivityNet Captions (Krishna et al., 2017), Charades-STA (Gao et al., 2017), and QVHighlights (Lei et al., 2021). We also perform evaluation on Video QA datasets such as ActivityNet-QA (Yu et al., 2019), MSRVTT-QA, and MSVD-QA (Xu et al., 2017) to evaluate Momentor in general question answering. Implementation details of Momentor can be found in Appendix B.
5.2 Action Segmentation
5.3 Dense Video Captioning
5.4 Temporal Grounding
5.5 Highlight Detection
Given a video and a description of the highlight activities within the video, highlight detection requires the model to output all the highlighted segments corresponding to the description. We evaluate Momentor on QVHighlights, and the results can be referenced in Table 3. From these results we can know: (1) Among all existing Video-LLMs, Momentor achieves state-of-the-art performance on highlight detection. (2) Thanks to the multi-event reasoning ability developed on the cross-segment tasks, Momentor can perceive the overall video semantics from a global perspective and effectively comprehend the relationships between different events, which is a key factor in highlight detection.
@@ 5.5 Highlight Detection
5.6 Video QA
5.7 In-Depth Analysis
Ablation Studies.
We conduct ablation experiments to assess the effectiveness of each component. The experiments are conducted under the following settings: (1) w/o continuous interpolation: We still use temporal tokens to express temporal positions, but without integrating the continuous interpolation mechanism. (2) w/o neighboring token propagation: We use the continuous temporal token space for temporal positioning, but without applying the neighboring token propagation mechanism when training. (3) w/o grounded event-sequence modeling: After modality alignment, we proceed directly to instruction fine-tuning without grounded event-sequence modeling. (4) w/o cross-segment tasks: We remove all instructions from cross-segment tasks and use only single-segment tasks for fine-tuning. We train Momentor with these settings and evaluate performances on ActivityNet Captions (temporal grounding and dense video captioning), Breakfast (action segmentation) and QVHighlights (highlight detection). The results of the ablation experiments can be referenced in Table 5.
@@ 5.7 In-Depth Analysis
Overall, removing any one of these components results in a decrease in the model’s overall performance. From Table 5, we can analyze the impact of removing different components on model performance separately. After removing the continuous interpolation mechanism, due to the quantization error, Momentor experiences a minor decline in localization-related metrics across all tasks, while the caption quality-related metrics are not significantly affected. Removing the neighboring token propagation mechanism leads to a performance drop in all metrics. Without neighboring token propagation, the temporal tokens are updated as multiple unrelated tokens rather than as an ordered sequence, which undermines the temporal representation and modeling. Visualizations of the temporal tokens (Section 8) also confirm this observation. Removing grounded event-sequence modeling leads to a significant performance decline in dense prediction tasks like dense video captioning and action segmentation, which indicates that grounded event-sequence modeling plays an important role in sequential semantics comprehension. The removal of cross-segment tasks has minimal impact on the performance of temporal grounding, as it does not involve cross-segment understanding. Performance on other tasks generally decreases, as both dense video captioning and action segmentation involve comprehension of multiple segments, and highlight detection also requires the model to distinguish between highlight segments and background segments.
@@
Figure 5: Dataset validation.Act.: ActivityNet.Break.: Breakfast.QV.: QVHighlights.
Figure 5: Dataset validation.Act.: ActivityNet.Break.: Breakfast.QV.: QVHighlights.
@@
Since temporal tokens are used to represent uniformly distributed temporal positions, we expect them to exhibit continuity in their embeddings. We use PCA (Abdi & Williams, 2010) and t-SNE (Van der Maaten & Hinton, 2008) to reduce the dimensionality of temporal tokens of Momentor and time tokens of Vid2Seq (Yang et al., 2023) into 1D and 2D for visualization. To validate the effectiveness of neighboring token propagation, we also visualize the temporal tokens trained without neighboring token propagation. For a fair comparison, we set the random state of t-SNE fixed to be 0. For the 1D reductions, we use the token indices as the x-axis and the reduced values as the y-axis; for the 2D reductions, we directly use the reduced values as coordinates. We employ a gradient color scheme, where the color of the data points will change progressively with the token index, as shown in Figure 8. It is evident that with neighboring token propagation, the embeddings of temporal tokens in Momentor are significantly more continuous. In contrast, embeddings of temporal tokens without neighboring token propagation and time tokens of Vid2Seq exhibit much less continuity, as their correlation can only be learned indirectly and inefficiently.
Figure 7: Analysis on special cases.
@@ Appendix D Prompts
Hello, I want you to act as a comprehensive video captioner. You will receive a list of frame-by-frame descriptions extracted from one video. Since some of these descriptions might be noisy, you should comprehend the major content of the video, assess the correctness of different pieces of information, and filter out erroneous, repetitive, noisy, or irrelevant details. After receiving and analyzing all the descriptions, please generate a comprehensive caption that effectively summarizes the events taking place in the video. Below are the information extracted from the video: {descriptions} There are some requirements that you should follow: 1. Your comprehensive video caption should be comprehensive, concise, informative, and LESS than 20 words. 2. You should output ONLY THE COMPREHENSIVE VIDEO CAPTION, and NO OTHER CONTENTS should be printed. 3. Your comprehensive video caption MUST NOT mention the concept of ”frame” or ”video”. Now please print out your comprehensive video caption. The comprehensive video caption:
@@
Generate a concise dialogue with factual questions and their answers based on the following video segment caption: {segment_caption} The answers should be directly inferred from the provided segment caption. Keep the questions and answers brief, with no more than 20 words each. Using ”User” and ”Assistant” as roles for questions and answers, respectively. Answer as if the ”Assistant” can directly watch the video. Speak as a friendly and helpful assistant.
@@
Generate a concise dialogue about the {instance_class} with factual questions and their answers based on the following video segment caption: {segment_caption} The answers should be directly inferred from the provided segment caption. Keep the questions and answers brief, with no more than 20 words each. Using ”User” and ”Assistant” as roles for questions and answers, respectively. Answer as if the ”Assistant” can directly watch the video. Speak as a friendly and helpful assistant.
@@
Hello, and I’d like you to act as a question generator. You will receive a sentence describing a clip in one video, and your task is to generate two questions with hypothetical scenario contexts to test if a deep learning model can retrieve the given clip based on the provided scenario information by asking about ”what scene would you see” and ”which clip should I watch” under that circumstance. Below are a few examples: <Example 1> [Clip Content] A elderly man is giving a speech in front of a blackboard. [Question] 1. Suppose you are a college student and you are in class one morning. Which clip might demonstrate the scene in front of you at this time? 2. If I want to know how older generations teach classes, which clip should I watch? <Example 2> [Clip Content] A young woman is seen standing in a room and dancing around. [Question] 1. You are a dance instructor. You are coaching your students in preparation for the next dance competition. What might you see at this moment? Please find the clip which might show this scene. 2. I want to learn to dance. Could you please tell me which clip of this video I should watch? <Example 3> [Clip Content] A dog in socks walks slowly out onto the floor as a lady films him. [Question] 1. A female animal behavior researcher is studying the walking patterns of dogs when their feet tend to slip in your laboratory. Which clip in the video might you see at this point? 2. I’m feeling anxious and need to watch some funny animal videos to relax. Could you please help me find such a clip in the video? Now given the following clip content sentence, please generate two questions with hypothetical scenario contexts to test if a deep learning model can retrieve the given clip based on the provided scenario information by asking about ”what scene would you see” and ”which clip should I watch” under that circumstance. [Clip Content] {content}
@@
You are a visual assistant. Given several video clip descriptions, you are tasked to generate a concise factual question and its answer by combining information from all the video clip descriptions. Note that the {instance_class} of all video clip descriptions in an input are the same, namely the video clip descriptions could describe the {instance_class} at different time points. The answer should be directly inferred from the provided sentence. Keep the question and answer brief. Using ”User” and ”Assistant” as roles for questions and answers, respectively. Speak as a friendly and helpful assistant. Below are some examples: <Example 1> [Input] 15.50s-30.75s : A group of children play in a park, running around and laughing. 45.20s-58.90s : A dog chases a frisbee, jumps to catch it, and returns it to its owner. [Output] User: What activities do the children and the dog engage in during the given video clips? Assistant: The children play in a park, running and laughing, while the dog chases a frisbee, jumps to catch it, and returns it to its owner. <Example 2> [Input] 10.50s-30.88s : A chef in a white apron chops vegetables on a wooden cutting board. 50.20s-63.40s : A close-up of a sizzling steak on a hot grill. 80.16s-92.74s : A chef takes freshly baked bread out of the oven and places it on a cooling rack. [Output] User: What cooking activities can be observed in the video? Assistant: The video shows a chef chopping vegetables, frying steak on a hot grill, and taking freshly baked bread out of the oven. Now given the following video clip descriptions, please generate a question-answer pair as it is in the examples. Note that the {instance_class} of all video clip descriptions in an input are the same, namely the video clip descriptions may show the events or actions surrounding the {instance_class} at different time points. The answer should be directly inferred from the provided sentence. Keep the question and answer brief. Using ”User” and ”Assistant” as roles for questions and answers, respectively. Speak as a friendly and helpful assistant. [Input] {segment_caption}
@@
Hello, I want you to act as a comprehensive video captioner. You will receive a list of clip-by-clip descriptions extracted from one video. Since some of these descriptions might be noisy, you should comprehend the major content of the video, assess the correctness of different pieces of information, and filter out erroneous, repetitive, noisy, or irrelevant details. After receiving and analyzing all these descriptions, please generate a comprehensive caption that effectively summarizes the events taking place in the video about the {instance_class}. Below are the clip descriptions extracted from the video: {descriptions} There are some requirements that you should follow: 1. Your comprehensive video caption should be comprehensive, concise, informative, and LESS than 20 words. 2. You should output ONLY THE COMPREHENSIVE VIDEO CAPTION, and NO OTHER CONTENTS should be printed. 3. Your comprehensive video caption MUST NOT mention the concept of ”frame” or ”video”. Now please print out your comprehensive video caption. The comprehensive video caption:

### v2

@@ 1 Introduction
Inspired by the success of ChatGPT (OpenAI, 2022), numerous studies across various fields are attempting to integrate Large Language Models (LLMs) with their domain-specific tasks, seeking to bring innovation to these fields. For example, Video Large Language Models (Video-LLMs) such as VideoChat (Li et al., 2023d) and Video-ChatGPT (Maaz et al., 2023) adapt LLM to video modality, striving to merge the understanding, reasoning and interactive skills of LLM with video perception. They typically sample multiple frames from the video, use an image encoder to encode these frames separately, and employ a projection layer (e.g. a linear layer or Q-Former (Li et al., 2023a)) to adapt the visual features to the feature space of an open-source LLM ((Touvron et al., 2023), (Chiang et al., 2023)). By training on video-level captioning and QA tasks, they establish coarse-grained multimodal feature alignment and acquire the capability of instruction following.
@@
We conduct extensive experiments with our proposed Momentor. The results indicate that our Momentor outperforms previous Video-LLMs in multiple tasks involving precise temporal position, such as temporal grounding, dense captioning, action segmentation, and highlight moment retrieval. Momentor demonstrates advanced proficiency in temporal perception. It can provide appropriate responses based on user-indicated segments as well as quickly locate target segments that meet user requirements.
@@ 2 Related Work
2.1 Vision and Language Understanding
With the rise of deep learning methods in the fields of computer vision and natural language processing, many efforts have been made to explore more complex multimodal understanding of vision and language. For example, tasks such as image and video-based QA, captioning and retrieval have been extensively discussed and explored by many existing studies (Antol et al., 2015; Vinyals et al., 2015; Faghri et al., 2017; Pan et al., 2023; Tapaswi et al., 2016; Venugopalan et al., 2015; Dong et al., 2021). Inspired by the success of the pre-training paradigm in natural language processing and computer vision, many works (Radford et al., 2021; Li et al., 2023a, 2022a; Sun et al., 2019) propose multimodal pre-trained models with excellent generalization by pre-training on a large amount of image-text or video-text pairs.
@@ 2.2 Temporally Grounded Video Understanding
Fine-grained video understanding tasks usually demand the model to view a video as a series of interconnected events and comprehend or locate them in a temporally grounded manner. For instance, action segmentation (Singh et al., 2016; Du et al., 2022; Behrmann et al., 2022) requires the model to temporally split the video and output the action label for each segment; temporal grounding (Gao et al., 2017; Zhang et al., 2020b, a; Li et al., 2022b, 2023c) demands the model to identify the start and end timestamps of the video segment corresponding to a given natural language query; highlight moment retrieval (Lei et al., 2021; Lin et al., 2023) requires the model to find out the central event in a video from a natural language description and pinpoint all related segments; dense video captioning (Alwassel et al., 2021; Yang et al., 2023a) requires the model to list out all events contained in a video along with their start and end timestamps. Previous methods typically train a task-specific model for each task, whereas we aim to design a unified Video-LLM that can solve these tasks in a zero-shot manner.
2.3 Multimodal Large Language Models
Many efforts have been made to transfer the task-handling capability of Large Language Models (LLMs) to the vision modality, enabling them to complete various tasks based on image content in accordance with user instructions (Liu et al., 2024; Li et al., 2023b; Pan et al., 2024; Zhu et al., 2023; Ge et al., 2024; Gao et al., 2024; Zhang et al., 2024). Several models (Li et al., 2023d; Maaz et al., 2023; Zhang et al., 2023; Luo et al., 2023; Huang et al., 2023; Ren et al., 2023) also incorporate temporal information aggregation module with LLM, in order that they can understand video content. Despite being effective in captioning or QA on short videos, the lack of fine-grained temporal modeling in these models prevents them from understanding or locating specific segments in long videos. In contrast, Momentor employs a Temporal Perception Module that integrates a continuous temporal token space for precise temporal positioning and modeling.
@@ 3.2 Temporal Perception Module (TPM)
Since each sampled frame is encoded and projected separately, their features do not contain the corresponding temporal position information. After constructing a continuous temporal token space and applying the neighboring token propagation, now we can actually obtain temporal embeddings corresponding to any timestamp, which contain precise temporal position information and possess the valuable property of temporal continuity. Therefore, we obtain the temporal embeddings at the positions of the sampled frames and directly add them to the projected frame features, as they share the same dimensionality, serving as a form of temporal position encoding to inject fine-grained temporal information.
@@ 3.3 Grounded Event-Sequence Modeling
After the stage of modality alignment, the model only learns the coarse-grained correspondence between visual and textual data. It still lacks fine-grained temporal awareness, so fine-tuning it directly on instruction data with precise timestamps can lead to slow convergence and ineffective event-sequence modeling. Therefore, we apply event-sequence decoding as an intermediary task that bridges the gap between low-level semantic alignment and high-level conceptual interaction. To be precise, given an untrimmed video as input, we require the model to output the event-sequence within it. We represent the $k$ th event as $E_{k}=[t_{start}^{k},t_{end}^{k},w_{1}^{k},...,w_{l_{k}}^{k}]$, where $t_{start}^{k}$, $t_{end}^{k}$ are the continuous temporal embeddings at the start and end of the $k$ th event, and $[w_{1}^{k},...,w_{l_{k}}^{k}]$ is a general caption composed of $l_{k}$ tokens for this event. The timestamps and general captions of each event in the event-sequence can be conveniently obtained during our instruction generation process without additional calculation (Section 4.2). We concatenate all the events to formulate the event-sequence $T_{E}=\{E_{i}\}_{i=1}^{N_{E}}$, where $N_{E}$ is the number of events in the untrimmed video. We apply a language modeling loss for event-sequence decoding:
@@ 4.1 Structured Information Extraction
Model Temporal Grounding Highlight Moment Retrieval
@@
Table 3: Comparison with existing Video-LLMs on temporal grounding and highlight moment retrieval
@@
Based on the result of instance tracking and event segmentation, we construct an Instance-Event Matrix, where each row represents an instance track (the video itself also counts as a track), and each column represents an event. The instance-event matrix shares certain similarities with video scene graphs (Shang et al., 2017; Yang et al., 2023b) as both involve instance behaviour tracking and structured semantic representation, but the instance-event matrix places greater emphasis on modeling the complex associations between events. We traverse the matrix and utilize several multimodal pre-trained models to extract visual clues such as scenes, instances, actions and attributes from each track. With the structured information organized in instance-event matrix, we can quickly generate instruction data that includes various spatio-temporal associations.
Figure 4: Distribution of different tasks in Moment-10M.
@@ 4.2 Instruction Generation
We feed the information in the instance-event matrix into Vicuna (Chiang et al., 2023), an open-source text-based LLM, to generate instruction data. We design various types of instruction-following tasks to comprehensively train and evaluate Video-LLMs. We incorporate 5 tasks focusing on single segment understanding as well as 3 tasks that involve reasoning across multiple segments, as shown in Figure 3. We utilize various prompts to guide Vicuna in generating instruction data for different tasks. Data from all 8 task types are used for instruction fine-tuning, while segment captioning data organized chronologically will be utilized for Grounded Event Sequence Modeling (Section 3.3). Detailed task descriptions and prompts can be found in Appendix C and D. We select a substantial number of videos from YTTemporal-1B (Zellers et al., 2022) to build Moment-10M. Figure 4 shows the distribution of each type of instruction data in Moment-10M. As shown in Table 1, Moment-10M comprises 10 million instruction data over 1.5 million segments and 451.5 thousand instance tracks. On average, each video contains 22.7 segments, which reflects the complexity of the event-sequences in the videos. We fine-tune Momentor on Moment-10M, enabling it to perform segment-level reasoning and localization.
@@ 5.1 Experiment Setup
To comprehensively evaluate Momentor in fine-grained understanding and precise localization, we assess it in a zero-shot setting across four tasks, i.e., action segmentation, dense video captioning, temporal grounding, and highlight moment retrieval, using datasets such as Breakfast (Kuehne et al., 2014), 50 Salads (Stein & McKenna, 2013), ActivityNet Captions (Krishna et al., 2017), Charades-STA (Gao et al., 2017), and QVHighlights (Lei et al., 2021). We also perform evaluation on Video QA datasets such as ActivityNet-QA (Yu et al., 2019), MSRVTT-QA, and MSVD-QA (Xu et al., 2017) to evaluate Momentor in general question answering. Implementation details of Momentor can be found in Appendix B.
@@
5.2 Action Segmentation
5.3 Dense Video Captioning
5.4 Temporal Grounding
5.5 Highlight Moment Retrieval
Given a video and a description of the highlight activities within the video, highlight moment retrieval requires the model to locate all the highlighted segments corresponding to the description. We evaluate Momentor on QVHighlights, and the results can be referenced in Table 3. From these results we can know: (1) Among all existing Video-LLMs, Momentor achieves state-of-the-art performance on highlight moment retrieval. (2) Thanks to the multi-event reasoning ability developed on the cross-segment tasks, Momentor can perceive the overall video semantics from a global perspective and effectively comprehend the relationships between different events, which is a key factor in highlight moment retrieval.
@@ 5.5 Highlight Moment Retrieval
5.6 Video QA
5.7 In-Depth Analysis
Ablation Studies.
We conduct ablation experiments to assess the effectiveness of each component. The experiments are conducted under the following settings: (1) w/o continuous interpolation: We still use temporal tokens to express temporal positions, but without integrating the continuous interpolation mechanism. (2) w/o neighboring token propagation: We use the continuous temporal token space for temporal positioning, but without applying the neighboring token propagation mechanism when training. (3) w/o grounded event-sequence modeling: After modality alignment, we proceed directly to instruction fine-tuning without grounded event-sequence modeling. (4) w/o cross-segment tasks: We remove all instructions from cross-segment tasks and use only single-segment tasks for fine-tuning. We train Momentor with these settings and evaluate performances on ActivityNet Captions (temporal grounding and dense video captioning), Breakfast (action segmentation) and QVHighlights (highlight moment retrieval). The results of the ablation experiments can be referenced in Table 5.
Overall, removing any one of these components results in a decrease in the model’s overall performance. From Table 5, we can analyze the impact of removing different components on model performance separately. After removing the continuous interpolation mechanism, due to the quantization error, Momentor experiences a minor decline in localization-related metrics across all tasks, while the caption quality-related metrics are not significantly affected. Removing the neighboring token propagation mechanism leads to a performance drop in all metrics. Without neighboring token propagation, the temporal tokens are updated as multiple unrelated tokens rather than as an ordered sequence, which undermines the temporal representation and modeling. Visualizations of the temporal tokens (Section 8) also confirm this observation. Removing grounded event-sequence modeling leads to a significant performance decline in dense prediction tasks like dense video captioning and action segmentation, which indicates that grounded event-sequence modeling plays an important role in sequential semantics comprehension. The removal of cross-segment tasks has minimal impact on the performance of temporal grounding, as it does not involve cross-segment understanding. Performance on other tasks generally decreases, as both dense video captioning and action segmentation involve comprehension of multiple segments, and highlight moment retrieval also requires the model to distinguish between highlight segments and background segments.
Figure 5: Dataset validation.
Act.: ActivityNet.
Break.: Breakfast.
QV.: QVHighlights.
@@ 5.7 In-Depth Analysis
Figure 7: Analysis on special cases.
@@
Since temporal tokens are used to represent uniformly distributed temporal positions, we expect them to exhibit continuity in their embeddings. We use PCA (Abdi & Williams, 2010) and t-SNE (Van der Maaten & Hinton, 2008) to reduce the dimensionality of temporal tokens of Momentor and time tokens of Vid2Seq (Yang et al., 2023a) into 1D and 2D for visualization. To validate the effectiveness of neighboring token propagation, we also visualize the temporal tokens trained without neighboring token propagation. For a fair comparison, we set the random state of t-SNE fixed to be 0. For the 1D reductions, we use the token indices as the x-axis and the reduced values as the y-axis; for the 2D reductions, we directly use the reduced values as coordinates. We employ a gradient color scheme, where the color of the data points will change progressively with the token index, as shown in Figure 8. It is evident that with neighboring token propagation, the embeddings of temporal tokens in Momentor are significantly more continuous. In contrast, embeddings of temporal tokens without neighboring token propagation and time tokens of Vid2Seq exhibit much less continuity, as their correlation can only be learned indirectly and inefficiently.
@@ 6 Conclusion
Acknowledgements
This work was supported by the Key Research and Development Projects in Zhejiang Province (No. 2024C01106), the NSFC (No. 62372341), the National Key Research and Development Project of China (2018AAA0101900), Ant Group, and Research funding from FinVolution Group.
@@ Appendix D Prompts
Segment Captioning Hello, I want you to act as a comprehensive video captioner. You will receive a list of frame-by-frame descriptions extracted from one video. Since some of these descriptions might be noisy, you should comprehend the major content of the video, assess the correctness of different pieces of information, and filter out erroneous, repetitive, noisy, or irrelevant details. After receiving and analyzing all the descriptions, please generate a comprehensive caption that effectively summarizes the events taking place in the video. Below are the information extracted from the video: {descriptions} There are some requirements that you should follow: 1. Your comprehensive video caption should be comprehensive, concise, informative, and LESS than 20 words. 2. You should output ONLY THE COMPREHENSIVE VIDEO CAPTION, and NO OTHER CONTENTS should be printed. 3. Your comprehensive video caption MUST NOT mention the concept of ”frame” or ”video”. Now please print out your comprehensive video caption. The comprehensive video caption:
@@
Segment QA Generate a concise dialogue with factual questions and their answers based on the following video segment caption: {segment_caption} The answers should be directly inferred from the provided segment caption. Keep the questions and answers brief, with no more than 20 words each. Using ”User” and ”Assistant” as roles for questions and answers, respectively. Answer as if the ”Assistant” can directly watch the video. Speak as a friendly and helpful assistant.
@@
Instance QA Generate a concise dialogue about the {instance_class} with factual questions and their answers based on the following video segment caption: {segment_caption} The answers should be directly inferred from the provided segment caption. Keep the questions and answers brief, with no more than 20 words each. Using ”User” and ”Assistant” as roles for questions and answers, respectively. Answer as if the ”Assistant” can directly watch the video. Speak as a friendly and helpful assistant.
@@
Inferential Segment Localization Hello, and I’d like you to act as a question generator. You will receive a sentence describing a clip in one video, and your task is to generate two questions with hypothetical scenario contexts to test if a deep learning model can retrieve the given clip based on the provided scenario information by asking about ”what scene would you see” and ”which clip should I watch” under that circumstance. Below are a few examples: <Example 1> [Clip Content] A elderly man is giving a speech in front of a blackboard. [Question] 1. Suppose you are a college student and you are in class one morning. Which clip might demonstrate the scene in front of you at this time? 2. If I want to know how older generations teach classes, which clip should I watch? <Example 2> [Clip Content] A young woman is seen standing in a room and dancing around. [Question] 1. You are a dance instructor. You are coaching your students in preparation for the next dance competition. What might you see at this moment? Please find the clip which might show this scene. 2. I want to learn to dance. Could you please tell me which clip of this video I should watch? <Example 3> [Clip Content] A dog in socks walks slowly out onto the floor as a lady films him. [Question] 1. A female animal behavior researcher is studying the walking patterns of dogs when their feet tend to slip in your laboratory. Which clip in the video might you see at this point? 2. I’m feeling anxious and need to watch some funny animal videos to relax. Could you please help me find such a clip in the video? Now given the following clip content sentence, please generate two questions with hypothetical scenario contexts to test if a deep learning model can retrieve the given clip based on the provided scenario information by asking about ”what scene would you see” and ”which clip should I watch” under that circumstance. [Clip Content] {content}
@@
Cross Segment QA You are a visual assistant. Given several video clip descriptions, you are tasked to generate a concise factual question and its answer by combining information from all the video clip descriptions. Note that the {instance_class} of all video clip descriptions in an input are the same, namely the video clip descriptions could describe the {instance_class} at different time points. The answer should be directly inferred from the provided sentence. Keep the question and answer brief. Using ”User” and ”Assistant” as roles for questions and answers, respectively. Speak as a friendly and helpful assistant. Below are some examples: <Example 1> [Input] 15.50s-30.75s : A group of children play in a park, running around and laughing. 45.20s-58.90s : A dog chases a frisbee, jumps to catch it, and returns it to its owner. [Output] User: What activities do the children and the dog engage in during the given video clips? Assistant: The children play in a park, running and laughing, while the dog chases a frisbee, jumps to catch it, and returns it to its owner. <Example 2> [Input] 10.50s-30.88s : A chef in a white apron chops vegetables on a wooden cutting board. 50.20s-63.40s : A close-up of a sizzling steak on a hot grill. 80.16s-92.74s : A chef takes freshly baked bread out of the oven and places it on a cooling rack. [Output] User: What cooking activities can be observed in the video? Assistant: The video shows a chef chopping vegetables, frying steak on a hot grill, and taking freshly baked bread out of the oven. Now given the following video clip descriptions, please generate a question-answer pair as it is in the examples. Note that the {instance_class} of all video clip descriptions in an input are the same, namely the video clip descriptions may show the events or actions surrounding the {instance_class} at different time points. The answer should be directly inferred from the provided sentence. Keep the question and answer brief. Using ”User” and ”Assistant” as roles for questions and answers, respectively. Speak as a friendly and helpful assistant. [Input] {segment_caption}
@@
Instance Activity Summarizing Hello, I want you to act as a comprehensive video captioner. You will receive a list of clip-by-clip descriptions extracted from one video. Since some of these descriptions might be noisy, you should comprehend the major content of the video, assess the correctness of different pieces of information, and filter out erroneous, repetitive, noisy, or irrelevant details. After receiving and analyzing all these descriptions, please generate a comprehensive caption that effectively summarizes the events taking place in the video about the {instance_class}. Below are the clip descriptions extracted from the video: {descriptions} There are some requirements that you should follow: 1. Your comprehensive video caption should be comprehensive, concise, informative, and LESS than 20 words. 2. You should output ONLY THE COMPREHENSIVE VIDEO CAPTION, and NO OTHER CONTENTS should be printed. 3. Your comprehensive video caption MUST NOT mention the concept of ”frame” or ”video”. Now please print out your comprehensive video caption. The comprehensive video caption: